(i) An isolated cubical conductor has an excess charge of 5.30 μC placed on its surface. Inside the conductor is a cavity, within which is a point charge of 4.10 μC. How many excess electrons are on the wall of the cavity?

(ii) Calculate the total electric flux that would pass through an irregularly shaped Gaussian surface drawn around the conductor in the previous problem.
Answer #1

(i) excess electrons = charge / charge on 1 electron

= 5.3*10^-6 / (1.6*10^-19)

= 3.3125*10^13

(ii) total electric flux = charge enclosed / epsilon  

= q/o

= (5.3 + 4.1) *10^-6 / 8.85*10^-12

= 1.062*10^6 Nm^2 / C